OpenStudy (anonymous):

x+6=sqaure root 18+3x+x^2

OpenStudy (anonymous):

\[x+6=\sqrt{18+3x+x ^{2}}\]

OpenStudy (anonymous):

solve the equation for x. if necessary round your answer to two decimal places.

OpenStudy (phi):

I would square both sides

OpenStudy (anonymous):

Rearrange the equation \[\sqrt{x^2 +3x+18} = x+6\] Square both sides\[x^2 +3x+18 = (x+6)^2\]\[x^2 +3x+18 = x^2 + 12x + 36\] Can you do the rest?

OpenStudy (anonymous):

i just put them into factors right?

OpenStudy (phi):

in this case, first combine *like terms*

OpenStudy (anonymous):

Like what "phi" said, combine the like terms. Here, the x squared cancel each other out. \[x^2 - x^2 = 0\]

OpenStudy (anonymous):

ok one moment

OpenStudy (anonymous):

x=-2 ?

OpenStudy (anonymous):

Correct!

OpenStudy (phi):

yes, but whenever you get this type of problem check that the answer works in the original equation (sometimes it won't!)

OpenStudy (anonymous):

ok, thanks guys!

OpenStudy (anonymous):

No problem :)